# Autonomous Dissection in Robotic Cholecystectomy

### v1

@@ Abstract
Robotic surgery offers enhanced precision and adaptability, paving the way for automation in surgical interventions. Cholecystectomy, the gallbladder removal, is particularly well-suited for automation due to its standardized procedural steps and distinct anatomical boundaries. A key challenge in automating this procedure is dissecting with accuracy and adaptability. This paper presents a vision-based autonomous robotic dissection architecture that integrates real-time segmentation, keypoint detection, grasping and stretching the gallbladder with the left arm, and dissecting with the other. We introduce an improved segmentation dataset based on videos of robotic cholecystectomy performed by various surgeons, incorporating a new “liver bed” class to enhance boundary tracking after multiple rounds of dissection. Our system employs state-of-the-art segmentation models and an adaptive boundary extraction method that maintains accuracy despite tissue deformations and visual variations. Moreover, we implemented an automated grasping and pulling strategy to optimize tissue tension before dissection upon our previous work. Ex vivo evaluations on porcine livers demonstrate that our framework significantly improves dissection precision and consistency, marking a step toward fully autonomous robotic cholecystectomy.
@@ I Introduction
Robotic-assisted surgery (RAS) continues revolutionizing modern medicine by enabling higher precision, reduced patient trauma, and faster recovery than traditional open or laparoscopic procedures. This favorable setting has spurred research into automating various surgical tasks such as tissue ablation [1], blunt dissection [2, 3], and suturing [4, 5, 6, 7, 8, 9]. For instance, the work in [10] demonstrated an algorithm for automated suturing based on visual tracking. At the same time, point cloud-based methods have been explored to model soft tissue deformations [2], estimate applied forces [11, 12], and for surgical planning [13].
Most prior approaches to task automation in RAS rely on endoscopic imaging, which poses challenges due to its two-dimensional representation and variable lighting conditions. Advanced segmentation and detection techniques are critical for overcoming these limitations. Recent breakthroughs brings new frameworks for subtask automations [14], with for example endoscopic servo control [15]. However, their methods often remain confined to pixel-level processing without establishing logical tissue connections [16], or realying on markers [17].
@@
To address these shortcomings, we present a new segmentation dataset based on the Comprehensive Robotic Cholecystectomy Dataset (CRCD) [19, 20], which includes robotic cholecystectomy procedures performed with the classic da Vinci system. In addition, we introduce a new class for tissue segmentation that corresponds to the region where the gallbladder used to be attached to the liver (liver bed) so the model can track the boundary even after multiple rounds of dissection. Furthermore, we train several state-of-the-art models for tissue instance segmentation, such as the MaskDINO [21] and YOLO11 [22], and compare the performance with the previous Detectron2 [23] model. We also develop an online boundary extraction method and an improved method for following the boundary. Finally, we propose a novel method to automatically grasp and pull the tissue before dissection. Together, the ability to track the boundary online and grasp the gallbladder represents a major step towards making our automated procedure much closer to how surgeons perform dissection. The effectiveness of our framework is successfully evaluated ex-vivo using porcine liver.
@@ II-A Revised Custom Segmentation Dataset
To address these challenges, we enriched the segmentation dataset by introducing a new label, liver bed, which represents the liver surface where the gallbladder was previously attached (Fig. 1). We used the Segment Anything Model 2 (SAM2)[24] to annotate videos from the Comprehensive Robotic Cholecystectomy Dataset (CRCD) [19, 20], a large-scale ex vivo robotic cholecystectomy dataset featuring procedures performed by multiple surgeons. This dataset closely aligns with our intended application and captures the dynamic tissue variations during surgery. SAM2 segments objects in single images or sequential video frames using prompts such as pixel-based cues or bounding boxes. However, SAM2 does not inherently classify the segmented regions, so we manually labeled each tissue category to create a structured training dataset. This allowed us to fine-tune state-of-the-art instance segmentation models, including YOLO11, for improved tissue recognition.
@@
Data Type Categories Train instances Test instances
Segmentation Pig Liver 25988 8690
Pig Gallbladder 25988 8690
Liver Bed 21660 7261
Keypoints FBF 5476 1372
PCH 7320 1831
@@ II-B Keypoints
We also expanded the keypoint dataset to have stronger keypoint tracking of the instruments. The annotations for surgical instruments, including Fenestrated Bipolar Forceps (FBF) and Permanent Cautery Hook (PCH), were manually performed using the COCO annotator [25]. Moreover, we used a different keypoint structure for each instrument to ensure robustness against common transformations. Fig. 2 illustrates the keypoint locations of the FBF and PCH. These keypoints are strategically placed on instrument regions with distinct colors and edges to optimize detection accuracy. Table I details the number of annotated instances for the extended instrument keypoint dataset. Both the segmentation and keypoint dataset follow Microsoft’s COCO format [26], ensuring compatibility with diverse object detection models.
@@ III-A Hardware Setup
Figure 3: The experimental setup: a porcine liver with the gallbladder attached is placed on a tray facing the ceiling. The endoscopic view and the generated point cloud from the stereo endoscope with the extracted features are shown in the top right corner.
Our system utilizes a first-generation da Vinci with the dVRK [27] and a Si model endoscope for enhanced image quality and reduced noise. To control the monopolar current from the electrosurgical unit (ESU), we interfaced the energy cables and console pedals with an Arduino UNO [20]. Additionally, we calibrated the stereo endoscope to generate 3D point clouds and performed a custom robot arm calibration to minimize dVRK errors [18] (Fig. 3). Moreover, we upgraded the implementation platform to ROS2 [28].
@@ III-B Image/Point Cloud Post-Processing
We applied the following image-processing steps to design an online version of tracking the boundary of interest and to align the FBF and PCH with the tissues for better performance. First, we used the classical skeletonization algorithm [29], which iteratively sweeps a fixed-size window over the binary mask of the object, removing pixels at each iteration until the image stops changing. Since the gallbladder has an oval-like shape, the 3D skeleton points derived from the point cloud effectively represent its medial surface (yellow points in the figures). The boundary of interest (purple points in the figures) is the interface between the liver and the gallbladder and corresponds to the points on the right side of the skeleton. By applying Principal Component Analysis (PCA) [30] to the adjacent and skeleton points, we extract the principal axis that points to the right side of the skeleton, aiding in precise boundary alignment.
@@ III-C Grasping
We divided the grasping task into three steps: align, grasp, and pull. First, we align the pose of the FBF to be parallel to the surface of the gallbladder of interest, which is the region between the boundary and skeleton points obtained from the post-processing step. We apply PCA to the points surrounding this surface to determine the alignment pose. We realigned the principal axes of the gallbladder surface when necessary to ensure that the secondary axis extends from the skeleton to the boundary points and the normal axis of the surface is oriented perpendicular to the gallbladder surface, directed along the viewing direction of the endoscope. This alignment guarantees consistency in orientation across different viewpoints. The resulting three principal axes approximate a local frame attached to the right-side surface of the gallbladder. The FBF frame is oriented based on this local frame (Fig. 4). Once aligned, the FBF moves along its negative x-axis, facing through the gallbladder surface, to reach the grasping point between the center of the skeleton and the boundary, offset by a certain distance along the grasping direction. After grasping, the FBF pulls the gallbladder along its negative z-axis, tangential to the right-side surface of the gallbladder, until the boundary becomes nearly linear, where the deviation of the boundary is below $0.5mm$ (Fig. 5).
@@ III-D Dissection
Similar to grasping, we first align the PCH to the gallbladder’s surface before performing the dissection. The alignment method is identical to grasping, except the frame axes are mapped based on the PCH’s pose rather than the FBF’s (Fig. 4). Unlike our previous approach, where the PCH followed a fixed trajectory, we now extract a target point from the online boundary. Initially, this target point is the first point of the current boundary, where boundary points are ordered clockwise from the gallbladder’s center. Once the PCH reaches the target point, the algorithm will search for the next one within the new boundary, selecting the point with the maximum distance—up to $1cm$—from its current position while maintaining alignment with its movement direction. The round of the dissection will terminate when finding a target point sufficiently far (over $1mm$) from the current PCH location is impossible.
@@ IV-C Procedure Performance
Fig. 4 and Fig. 5 illustrate the stages of gallbladder grasping. The success of the grasping phase does not heavily depend on the models used; instead, the primary objective is to ensure that the boundary is sufficiently stretched for dissection and that the FBF remains stable throughout the procedure. Consequently, our focus is primarily on evaluating the dissection performance after the gallbladder has been adequately stretched, allowing for a direct comparison with results from our previous work. The FBF was not repositioned after the initial grasping to ensure consistency across trials. For Detectron2, we used DT2-seg and DT2-kpt to maintain the previous experimental setup, with both models fine-tuned on the new dataset. We then compared their performance with both our prior results and state-of-the-art models.
@@
Fig. 8 presents the recorded boundary points from one of the trials, comparing YOLO11l to DT2, demonstrating that the YOLO’s boundary remained stable over time. The results indicate that the dissection process did not significantly alter the boundary, even though boundary updates were performed in real-time. This validates the effectiveness of our grasping strategy in maintaining boundary stability during dissection, which was a key limitation in our previous study. We fitted a cubic spline to the recorded boundary points to quantify boundary consistency, computing the residual mean squared error (RMSE) between the points and the fitted curve, as shown in Table IV. Since the boundary points were fixed in our previous work, we excluded those results from the table. Additionally, we measured the total travel distance of the PCH during dissection and the time required to complete one full boundary dissection, including data from our prior study.
@@ V Conclusion
Furthermore, we transitioned from offline to online boundary point extraction, improving the methodology to ensure robustness under challenging surgical conditions, including multiple dissections, tissue color changes over time, and gallbladder size and shape variations. Additionally, we aligned the surgical instruments to the gallbladder surface before each action, enhancing task performance. We also introduced an automated bimanual manipulation strategy that integrates grasping and pulling actions, which are crucial for maintaining optimal tissue tension and compensating for shifts caused by energy application. To validate our framework, we conducted ex-vivo evaluations on porcine livers, demonstrating improved surgical instrument localization and precise dissection, even after the tissue deformation.
@@
Hence, future work will focus on adding these different dissection techniques to our framework. Although the keypoint detection has improved from our previous work, it still suffered in certain poses or situations during the task. So we plan to further augment the dataset for the keypoints using the CRCD dataset, similar to how we expanded the segmentation dataset. We also aim to refine our control strategy to fully separate the gallbladder from the liver, accomplishing multiple rounds of pulling and dissecting the gallbladder autonomously and comparing the performance with the surgeons.

### v2

@@ Abstract
Robotic surgery offers enhanced precision and adaptability, paving the way for automation in surgical interventions. Cholecystectomy, the gallbladder removal, is particularly well-suited for automation due to its standardized procedural steps and distinct anatomical boundaries. A key challenge in automating this procedure is dissecting with accuracy and adaptability. This paper presents a vision-based autonomous robotic dissection architecture that integrates real-time segmentation, keypoint detection, grasping and stretching the gallbladder with the left arm, and dissecting with the other arm. We introduce an improved segmentation dataset based on videos of robotic cholecystectomy performed by various surgeons, incorporating a new “liver bed” class to enhance boundary tracking after multiple rounds of dissection. Our system employs state-of-the-art segmentation models and an adaptive boundary extraction method that maintains accuracy despite tissue deformations and visual variations. Moreover, we implemented an automated grasping and pulling strategy to optimize tissue tension before dissection upon our previous work. Ex vivo evaluations on porcine livers demonstrate that our framework significantly improves dissection precision and consistency, marking a step toward fully autonomous robotic cholecystectomy.
@@ I Introduction
Robotic-assisted surgery (RAS) continues to revolutionize modern medicine by enabling higher precision, reduced patient trauma, and faster recovery than traditional open or laparoscopic procedures. This favorable setting has spurred research into automating various surgical tasks such as tissue ablation [1], blunt dissection [2, 3], and suturing [4, 5, 6, 7, 8, 9]. For instance, the work in [10] demonstrated an algorithm for automated suturing based on visual tracking. At the same time, point cloud-based methods have been explored to model soft tissue deformations [2], estimate applied forces [11, 12], and for surgical planning [13].
Most prior approaches to task automation in RAS rely on endoscopic imaging, which poses challenges due to its two-dimensional representation and variable lighting conditions. Advanced segmentation and detection techniques are critical for overcoming these limitations. Recent breakthroughs bring new frameworks for subtask automations [14], with, for example, endoscopic servo control [15]. However, their methods often remain confined to pixel-level processing without establishing logical tissue connections [16], or relying on markers [17].
@@
To address these shortcomings, we present a new segmentation dataset based on the Comprehensive Robotic Cholecystectomy Dataset (CRCD) [19], which includes robotic cholecystectomy procedures performed with the classic da Vinci system. In addition, we introduce a new class for tissue segmentation that corresponds to the region where the gallbladder used to be attached to the liver (liver bed), so the model can track the boundary even after multiple rounds of dissection. Furthermore, we train several state-of-the-art models for tissue instance segmentation, such as the MaskDINO [20] and YOLO11 [21], and compare the performance with the previous Detectron2 [22] model. We also develop an online boundary extraction method and an improved method for following the boundary. Finally, we propose a novel method to automatically grasp and pull the tissue before dissection. Together, the ability to track the boundary online and grasp the gallbladder represents a major step towards making our automated procedure much closer to how surgeons perform dissection. The effectiveness of our framework is successfully evaluated ex vivo using porcine liver.
@@ II-A Revised Custom Segmentation Dataset
To address these challenges, we enriched the segmentation dataset by introducing a new label, liver bed, which represents the liver surface where the gallbladder was previously attached (Fig. 1). We used the Segment Anything Model 2 (SAM2) [23] to annotate videos from the Comprehensive Robotic Cholecystectomy Dataset (CRCD) [19], a large-scale ex vivo robotic cholecystectomy dataset featuring procedures performed by multiple surgeons. This dataset closely aligns with our intended application and captures the dynamic tissue variations during surgery. SAM2 segments objects in single images or sequential video frames using prompts such as pixel-based cues or bounding boxes. However, SAM2 does not inherently classify the segmented regions, so we manually labeled each tissue category to create a structured training dataset. This allowed us to fine-tune state-of-the-art instance segmentation models, including YOLO11, for improved tissue recognition.
@@
Data Type Categories Previous [18] New
Train Test Train Test
Segmentation Pig Liver 1430 356 25988 8690
Pig Gallbladder 1429 359 25988 8690
Liver Bed - - 21660 7261
Keypoints FBF 471 130 5476 1372
PCH 1904 477 7320 1831
@@ II-B Keypoints
We also expanded the keypoint dataset to have stronger keypoint tracking of the instruments. The annotations for surgical instruments, including Fenestrated Bipolar Forceps (FBF) and Permanent Cautery Hook (PCH), were manually performed using the COCO annotator [24]. Moreover, we used a different keypoint structure for each instrument to ensure robustness against common transformations. Fig. 2 illustrates the keypoint locations of the FBF and PCH. These keypoints are strategically placed on instrument regions with distinct colors and edges to optimize detection accuracy. Table I details the number of annotated instances for the extended instrument keypoint dataset. Both the segmentation and keypoint datasets follow Microsoft’s COCO format [25], ensuring compatibility with diverse object detection models. The annotated datasets are included in the Expanded CRCD [26], which provides additional details and is publicly available.
@@
Figure 3: The overall system architecture: (a) Hardware setup: in addition to the dVRK, the ESU is included. Stereo endoscopic images are used to generate real-time 3D point clouds. (b) & (c) Outputs of the perception models: tissue instance segmentation and instrument keypoint detection. These results are used to compute the 3D dissection boundary and the 3D pose of each surgical instrument. (d) Grasping mechanism using the FBF (Section III-C). (e) Updated dissecting mechanism controlling the PCH (Section III-D).
@@ III-A Hardware Setup
Our system utilizes a first-generation da Vinci with the dVRK [27] and a Si model endoscope for enhanced image quality and reduced noise. To control the monopolar current from the electrosurgical unit (ESU), we interfaced the energy cables and console pedals with an Arduino UNO [26]. Additionally, we calibrated the stereo endoscope to generate 3D point clouds and performed a custom robot arm calibration to minimize dVRK errors [18] (Fig. 3(a)). Moreover, we upgraded the implementation platform to ROS2 [28].
@@ III-B Image/Point Cloud Post-Processing
We applied the following image-processing steps directly to the tissue segmentation masks provided from the models to design an online version of tracking the boundary of interest and to align the FBF and PCH with the tissues for better performance. First, we used the classical skeletonization algorithm [29], which iteratively sweeps a fixed-size window over the binary mask of the object, removing pixels at each iteration until the image stops changing. Since the gallbladder has an oval-like shape, the 3D skeleton points derived from the point cloud effectively represent its medial surface (yellow points in the figures). The boundary of interest (purple points in the figures) is the interface between the liver and the gallbladder and corresponds to the points on the right side of the skeleton. By applying Principal Component Analysis (PCA) [30] to the adjacent and skeleton points, we extract the principal axis that points to the right side of the skeleton, aiding in precise boundary alignment.
@@ III-C Grasping
We divided the grasping task into three steps: align, grasp, and pull. First, we align the pose of the FBF to be parallel to the surface of the gallbladder of interest, which is the region between the boundary and skeleton points obtained from the post-processing step. We apply PCA to the points surrounding this surface to determine the alignment pose. We realigned the principal axes of the gallbladder surface when necessary to ensure that the secondary axis extends from the skeleton to the boundary points and the normal axis of the surface is oriented perpendicular to the gallbladder surface, directed along the viewing direction of the endoscope. This alignment guarantees consistency in orientation across different viewpoints. The resulting three principal axes approximate a local frame attached to the right-side surface of the gallbladder. The FBF frame is oriented based on this local frame (Fig. 4). Once aligned, the FBF moves along its negative x-axis, facing through the gallbladder surface, to reach the grasping point between the center of the skeleton and the boundary, offset by a certain distance along the grasping direction. After grasping, the FBF pulls the gallbladder along its negative z-axis, tangential to the right-side surface of the gallbladder, until the boundary becomes nearly linear, where the deviation of the boundary is below $0.5mm$ (Fig. 5). This grasping mechanism stabilizes the gallbladder throughout the task, reducing the impact of anatomical variability and sudden tissue deformation.
@@ III-D Dissection
Once the grasping and pulling are accomplished, the FBF remains static. Similar to grasping, we first align the PCH to the gallbladder’s surface before performing the dissection. The alignment method is identical to grasping, except that the frame axes are mapped based on the PCH’s pose rather than the FBF’s (Fig. 4). Unlike our previous approach, where the PCH followed a fixed, preplanned trajectory, the dissection is now guided online by visual feedback. A target point is dynamically selected from the current tissue boundary, which is updated in real time to reflect changes in the gallbladder’s shape. Initially, the target point is the first in the boundary sequence, ordered clockwise from the center. As the PCH reaches each target, a new one is selected from the updated boundary, choosing the point furthest away (up to 1 cm) in the direction of motion, while maintaining proper alignment. The dissection round ends when no boundary point sufficiently distant (greater than 1 mm) from the current position can be found, ensuring the system adapts continuously to tissue deformation throughout the procedure.
@@ IV-C Procedure Performance
Fig. 4 and Fig. 5 illustrate the stages of gallbladder grasping. The success of the grasping phase does not heavily depend on the models used; instead, the primary objective is to ensure that the boundary is sufficiently stretched for dissection and that the FBF remains stable throughout the procedure. This evaluation also serves as an ablation analysis: our previous system [18] lacked grasping, alignment, and used older perception models, enabling us to isolate the impact of each updated module in the current framework. Consequently, our focus is primarily on evaluating the dissection performance after the gallbladder has been adequately stretched, allowing for a direct comparison with results from our previous work. The FBF was not repositioned after the initial grasping to ensure consistency across trials. For Detectron2, we used DT2-seg and DT2-kpt to maintain the previous experimental setup, with both models fine-tuned on the new dataset. We then compared their performance with both our prior results and state-of-the-art models.
@@
Fig. 8 presents the recorded boundary points from one of the trials, comparing YOLO11l to DT2, demonstrating that YOLO’s boundary remained stable over time. The results indicate that the dissection process did not significantly alter the boundary, even though boundary updates were performed in real-time. This validates the effectiveness of our grasping strategy in maintaining boundary stability during dissection, which was a key limitation in our previous study. We fitted a cubic spline (blue lines in Fig. 8) to the recorded boundary points to quantify boundary consistency, computing the residual mean squared error (RMSE) between the points and the fitted curve, as shown in Table IV. Since the boundary points were fixed in our previous work, we excluded those results from the table. Additionally, we measured the total travel distance of the PCH during dissection and the time required to complete one full boundary dissection, including data from our prior study. We do not explicitly report a task success rate or boundary coverage metric, as the dissection boundary is adaptively updated in real time to account for tissue deformation while performing the task.
@@ V Conclusion
Furthermore, we transitioned from offline to online boundary point extraction, improving the methodology to ensure robustness under challenging surgical conditions, including multiple dissections, tissue color changes over time, and gallbladder size and shape variations. Additionally, we aligned the surgical instruments to the gallbladder surface before each action, enhancing task performance. We also introduced an automated bimanual manipulation strategy that integrates grasping and pulling actions, which are crucial for maintaining optimal tissue tension and compensating for shifts caused by energy application. To validate our framework, we conducted ex vivo evaluations on porcine livers, demonstrating improved surgical instrument localization and precise dissection, even after tissue deformation.
@@
Although the keypoint detection has improved from our previous work, it still suffers in certain poses or situations during the task. To address this, we plan to further augment the dataset for keypoints using the CRCD dataset, similar to how we expanded the segmentation dataset. In parallel, we are developing a robustness strategy in which the endoscope actively follows the instrument tip, ensuring it remains in the center of the image to mitigate occlusion and boundary visibility issues during keypoint detection. We also aim to refine our control strategy to fully separate the gallbladder from the liver, accomplishing multiple rounds of pulling and dissecting the gallbladder autonomously and comparing the performance with the surgeons.